OpenStudy (anonymous):

If x-4 is 2 greater than y+1 then by how much is x+6 greater than y

OpenStudy (anonymous):

Start by stating "x-4 is 2 greater than y+1" as an equation: x - 4 = y + 1 + 2

OpenStudy (anonymous):

And i get x=y+7

OpenStudy (anonymous):

Good, so try to make the equation say x+6=something

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

Add 6 to both sides so that you can see what x+6 equals. If you know what x+6 equals, you will be able to see how much larger it is than y.